OpenStudy (anonymous):

Find the y-intercept of the exponential function f(x) = (4)^x - 6. A. (0, 0) B. (0, -1) C. (0, -5) D. (0, -6)

OpenStudy (anonymous):

@R00nnYBraiNsbiG

jimthompson5910 (jim_thompson5910):

plug in x = 0 and tell me what you get

OpenStudy (anonymous):

ok do u just want the answer or do u want 2 kno how to do it?

OpenStudy (anonymous):

answer, and help.

OpenStudy (anonymous):

jim I got 0

jimthompson5910 (jim_thompson5910):

f(x) = (4^x) - 6 f(0) = (4^0) - 6 f(0) = (1) - 6 ... anything, but 0, to the 0th power is 1 f(0) = ???

OpenStudy (anonymous):

well looks like he got this one

OpenStudy (anonymous):

D?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you have 1-6 not just -6

OpenStudy (anonymous):

ya but 1, -6 is not in my multiple choice

OpenStudy (radar):

No but when you have 1 -6 you get -5, now look again!

OpenStudy (radar):

X = 0 y = -5 Don't you see it?

OpenStudy (anonymous):

oh its C.

OpenStudy (radar):

Bingo !